OpenStudy (anonymous):

find all complex cube roots of -i. in other words find all complex solutions of x^3=-i

OpenStudy (kc_kennylau):

1. express -i in the form of cos(x)+isin(x)

OpenStudy (anonymous):

\[\pm \frac{ 1 }{ 2 }\left( \sqrt{3}-i \right)\]

OpenStudy (kc_kennylau):

The minus one is not correct